OpenStudy (anonymous):

to pay your monthly bills, you can either open a checking account or use and online banking service. a local bank service charges $3 per month and $.40 per check, while online services charges a flat fee of $9 per month If you pay about 12 bills a month which service should you choose? Justify your answer.

OpenStudy (mathstudent55):

The question is how many checks do you need to write per month to make the online banking service be a better deal? Let x = number of checks per month Local bank cost per month: 3 + 0.4x Online banking service cost per month: 9 3 + 0.4x = 9 Solve for x. If you write fewer than x checks per month, the local bank is a better deal. If you write x checks per month, they are the same. If you write more than x checks per month, the online banking service is a better deal.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

lets put it this way, put them into equations: the local bank: y=.4x+3 online bank: y=9x now plug in 12 for x and see which one would be better for banking. (which one is less?) :)

OpenStudy (anonymous):

show me please?

OpenStudy (anonymous):

all you have to do is put the 12 in for x and solve.. so it would be 4(12)+3 then you would do (4*12)+3 then do the same with the other one 9(12) and see which one is lower..

OpenStudy (anonymous):

so 48+3=108?

OpenStudy (anonymous):

you just do 48+3

OpenStudy (anonymous):

i got 108 and 51 is this right?

OpenStudy (anonymous):

yesh but you do not equal them up... so which one would you want to chose?

OpenStudy (anonymous):

51

OpenStudy (anonymous):

right so you would want to choose the local bank

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

medal? :)

OpenStudy (anonymous):

u r the best

OpenStudy (anonymous):

lol not really but thanks